# Closed-Form Power and Sample Size Calculations for Bayes Factors

### v1

@@ Abstract
Determining an appropriate sample size is a critical element of study design, and the method used to determine it should be consistent with the planned analysis. When the planned analysis involves Bayes factor hypothesis testing, the sample size is usually desired to ensure a sufficiently high probability of obtaining a Bayes factor indicating compelling evidence for a hypothesis, given that the hypothesis is true. In practice, Bayes factor sample size determination is typically performed using computationally intensive Monte Carlo simulation. Here, we summarize alternative approaches that enable sample size determination without simulation. We show how, under approximate normality assumptions, sample sizes can be determined numerically, and provide the R package bfpwr for this purpose. Additionally, we identify conditions under which sample sizes can even be determined in closed-form, resulting in novel, easy-to-use formulas that also help foster intuition, enable asymptotic analysis, and can also be used for hybrid Bayesian/likelihoodist design. Furthermore, we show how in our framework power and sample size can be computed without simulation for more complex analysis priors, such as Jeffreys-Zellner-Siow priors or nonlocal normal moment priors. Case studies from medicine and psychology illustrate how researchers can use our methods to design informative yet cost-efficient studies.
Keywords: Bayesian hypothesis testing, evidence, likelihood ratio, predictive power, study design
@@ 1 Introduction
A key aspect of study design is determining an appropriate sample size. Choosing a sample size that is too small may lead to inconclusive study results, while choosing a sample size that is too large may be unethical (e.g., for animal studies) or waste samples which could be of better use in other studies. Whether or not a certain sample size can ensure sufficiently conclusive results depends on the planned analysis. Therefore, the sample size calculation should be aligned with the planned analysis (Anderson and Kelley,, 2022), or in other words: ‘As ye shall analyse is as ye shall design(Julious,, 2023, p. 179).
@@
where $z_{q}$ is the $q\times 100\%$ quantile of the standard normal distribution, $\alpha$ is the level of the test, $1-\beta$ is the desired power, $\mu$ is the assumed mean difference, $\theta_{0}$ is the mean difference under the null hypothesis, and $\sigma^{2}$ is the variance of one observation (Matthews,, 2006, p. 34). There exist various refinements of (1), such as, adaptations to unequal randomization, special study designs (e.g., cross-over studies), or other data types (e.g., binary data), see, for example, Kieser, (2020) or Julious, (2023). For many analysis methods, however, no closed-form formula exist and iterative or simulation methods have to be used. Nevertheless, while typically only being an approximation, the formula (1) enables quick-and-dirty calculations that are often accurate enough for practical purposes. It also helps fostering intuition and is therefore useful, for example, in teaching of statistics or the study of asymptotics.
An alternative to frequentist hypothesis testing is Bayesian hypothesis testing. There are different flavors of Bayesian hypothesis testing, one of the most popular being the approach centered around the Bayes factor, which is the data-based updating factor of the prior to posterior odds of two competing hypotheses. Bayes factor approaches were pioneered by Jeffreys, (1939) and are now in use in various scientific domains such as medicine (Goodman,, 1999), psychology (Wagenmakers,, 2007), or physics (Trotta,, 2008). Bayes factor tests are conceptually different from frequentist tests in several ways. For example, they can quantify evidence in favor of a null hypothesis or they can incorporate external information via a prior distribution. For an overview of Bayes factors see e.g., Kass and Raftery, (1995); Held and Ott, (2018).
@@
The goal of this paper is therefore to investigate whether, under approximate normality assumptions similar to those underlying the formula (1), a sample size formula can be derived for a planned Bayes factor analysis. As we will show, the answer is affirmative under certain assumptions about the analysis prior (the prior distribution for the parameter used in the analysis) and the design prior (the prior distribution for the parameter used in the design). A distinction must be made between point priors and normal priors, both in the design and in the analysis. Point analysis priors lead to Bayes factors reducing to likelihood ratios, the analysis thereby corresponding with a frequentist ‘likelihoodist’ analysis (Edwards,, 1971; Royall,, 1997; Blume,, 2002; Strug,, 2018), while point design priors lead to ‘conditional power’, which corresponds to traditional frequentist power. In contrast, normal analysis and design priors can account for parameter uncertainty, producing Bayes factors that differ from likelihood ratios and ‘predictive power’ that differs from frequentist power (see e.g., O'Hagan et al.,, 2005; Micheloud and Held,, 2022; Grieve,, 2022, for conditional/predictive power related to posterior tail probability analyses).
Based on the point/normal prior distinction, we find that closed-form sample sizes are available for Bayes factors with point analysis priors (i.e., likelihood ratios) along with point or normal design priors, and for Bayes factors with local normal analysis and design priors (i.e., normal priors centered on the null value), see Table 1 for an overview of our results. In addition to our novel formulas, we summarize sample size determination for Bayes factors with normal priors based on numerical root-finding (which has done before, e.g., in Weiss,, 1997; Pawel et al.,, 2023), and show how it can be extended to more advanced prior distributions, such as normal moment priors and Jeffreys-Zellner-Siow priors. While root-finding approaches also do not produce closed-form sample sizes, computations are deterministic and usually faster than with simulation approaches. To facilitate reuse of our results, all methods are made available through our R package bfpwr.
@@
This paper is organized as follows: We begin by defining the type of Bayes factor underlying our sample size calculations (Section 2), followed by deriving its distribution when new data are generated under various design priors (Section 3). Combining these results, we derive several formulas for the sample size under different constellations of design and analysis priors (Section 4). Examples from medicine and psychology then illustrates how our formulas can be used by researchers in practice (Section 5). Section 6 illustrates possible extensions of our framework to other popular types of analysis priors, such as informed $t$ priors (Gronau et al.,, 2020) and nonlocal normal moment priors (Johnson and Rossell,, 2010). The paper ends with a closing discussion of our results and final remarks on limitations and extensions (Section 7). Our R package bfpwr that implements the developed methods is illustrated in Appendix A.
2 Assumed Bayes factor analysis
@@
To derive a sample size formula, we must first clarify how the future data will be analyzed. Denote by $\hat{\theta}$ the estimate of an unknown parameter $\theta$ that will result from the statistical analysis of $n$ effective future observations, each with unit variance $\sigma^{2}_{\scriptscriptstyle\hat{\theta}}$. Assume that the estimate (but not necessarily the underlying data) is approximately normally distributed around $\theta$ with variance $\sigma^{2}_{\scriptscriptstyle\hat{\theta}}/n$, i.e., $\hat{\theta}\mid\theta\sim\mathrm{N}(\theta,\sigma^{2}_{\scriptscriptstyle\hat{\theta}}/n)$. Table 2 shows common types of parameter estimates and the resulting interpretation of $n$ and the unit variance $\sigma^{2}_{\scriptscriptstyle\hat{\theta}}$. It is important to note that these are only approximations and they may be inadequate in certain situations, such as small sample sizes or rare events (Spiegelhalter et al.,, 2004). The frequentist sample size formula (1) can also be cast in this framework; it assumes continuous outcome data and a mean difference parameter (second row in Table 2). As such, the formula could be generalized to other settings by replacing $2\sigma^{2}$ in the numerator with other unit variances from Table 2, which would in turn change the interpretation of $n$ and $\hat{\theta}$.
@@
Table 2: Different types of parameter estimates $\hat{\theta}$ with approximate variance $\mathrm{Var}(\hat{\theta})=\sigma^{2}_{\scriptscriptstyle\hat{\theta}}/n$ and corresponding interpretation of sample size $n$ and unit variance $\sigma^{2}_{\scriptscriptstyle\hat{\theta}}$ (adapted from Chapter 2.4 in Spiegelhalter et al.,, 2004 and Chapter 1 in Grieve,, 2022). The variance of one continuous outcome observation is denoted by $\sigma^{2}$. Parameter estimates based on two groups assume an equal number of observations per group.
@@
Binary Arcsine square root difference Sample size per group 1/2
@@
A Bayes factor smaller than one ($\text{BF}_{01}<1$) indicates evidence for the alternative hypothesis $H_{1}$, while a Bayes factor greater than one ($\text{BF}_{01}>1$) indicates evidence for the null hypothesis $H_{0}$. The larger the deviation of the Bayes factor from one, the stronger the evidence. The Bayes factor (2) is implemented in our package in the function bf01.
The Bayes factor (2) has already appeared in the literature in one form or another (e.g., in Weiss,, 1997; De Santis,, 2004; Spiegelhalter et al.,, 2004; Dienes,, 2014; Bartoš and Wagenmakers,, 2023), with perhaps the first proposal of a Bayes factor based on an approximately normally distributed parameter estimate and its standard error dating back to Jeffreys, (1936), see also Wagenmakers, (2022) for some historical notes on Jeffreys’ approach. The Bayes factor (2) may be thought of as a ‘Bayesian $z$-test’, that is, a test of a normal mean based on an asymptotically normal statistic assuming that the variance of the statistic is known. Of course, the latter assumption is not true in most applications, but it makes the test widely applicable and is, for practical purposes, often close enough to Bayes factors based on the exact distribution of the data, which may or may not be available. Finally, the Bayes factor (2) can also be used with parameter estimates where a standard error is available but not of the form $\sigma_{\scriptscriptstyle\hat{\theta}}/\sqrt{n}$, e.g., a parameter estimate from a generalized linear model where the estimate is adjusted for covariates and the standard error is obtained numerically. In this case, $\sigma_{\scriptscriptstyle\hat{\theta}}/\sqrt{n}$ in (2) can be replaced by the observed standard error. However, as we will show now, assuming such a particular dependence on the sample size $n$ allows us to perform closed-form power and sample size calculations under certain additional assumptions.
@@ 3 Distribution and power function of the Bayes factor
Suppose now that we are interested in finding compelling evidence in favor of the alternative $H_{1}$ over the null hypothesis $H_{0}$ with a Bayes factor (2) smaller than some threshold $k<1$. For example, the evidence thresholds could be $k=1/3$ or $k=1/10$, the levels from Jeffreys, (1939) for ‘substantial’ and ‘strong’ evidence, respectively. To determine a sample size that ensures compelling evidence with a desired probability we need to know the distribution of the Bayes factor (2) for a given sample size.
Assume a so-called ‘design prior’ for the parameter $\theta$ that is used in the design of the study (O'Hagan and Stevens,, 2001; O'Hagan et al.,, 2005). This prior should represent the state of knowledge and uncertainty about $\theta$ at the design stage and does not necessarily have to correspond to the ‘analysis prior’ $\theta\mid H_{1}\sim\mathrm{N}(\mu,\tau^{2})$ used for the Bayes factor (2). In fact, the analysis prior is often set to a certain ‘default’ or ‘objective’ prior that is conventionally used in the field. Here, we will focus on normal design priors $\theta\sim\mathrm{N}(\mu_{d},\tau^{2}_{d})$, as they are flexible enough to specify varying degrees of uncertainty about the parameter, and at the same time mathematically convenient for obtaining closed-form solutions for power and, in some cases, sample size. Point priors, and as such classical sample size determination with an ‘assumed parameter’, then represent a special case of the normal prior where the standard deviation becomes infinitesimally small ($\tau_{d}=0$).
@@
$Z=\frac{1}{\sqrt{\tau^{2}_{d}+\sigma^{2}_{\scriptscriptstyle\hat{\theta}}/n}}\left\{\frac{\sigma^{2}_{\scriptscriptstyle\hat{\theta}}\log k}{n(\theta_{0}-\mu)}+\frac{\theta_{0}+\mu}{2}-\mu_{d}\right\},$
see Appendix B for details. In standard frequentist sample size determination, the power can typically be increased arbitrarily close to one by increasing the sample size. However, with the Bayes factor based on a point analysis prior, depending on the assumed design prior, one may not be able to approach a power of one with the power function (3) by increasing the sample size $n$. That is, the limiting power value is given by
@@
with $Z_{\mathrm{lim}}=(\theta_{0}+\mu-2\mu_{d})/(2\tau_{d})$. When the design prior is also point prior ($\tau_{d}=0$), $Z_{\mathrm{lim}}$ diverges and the limiting power (4) approaches one or zero, depending on whether the location of the design prior $\mu_{d}$ is closer to the alternative $\mu$ or to the null $\theta_{0}$. In case it is just in between the two ($\mu_{d}=(\theta_{0}+\mu)/2$), the limiting power approaches a half. On the other hand, for a normal design prior ($\tau_{d}>0$), the limiting power is bounded by a value in between (and not including) zero and one given by (4). The intuition behind these results is that for a normal design prior, there is always parameter uncertainty, even if the sample size becomes arbitrarily large, while for point design priors, the parameter uncertainty can be arbitrarily reduced by increasing the sample size. This parallels similar results on bounds for hybrid Bayesian/frequentist power (Spiegelhalter et al.,, 2004; Micheloud and Held,, 2022; Grieve,, 2022).
@@
The left plot in Figure 1 shows examples of two power functions related to the Bayes factor for a standardized mean difference parameter with point analysis prior. The power is computed under either a point (red curve) or a normal design prior (blue curve). We see that increasing the sample size $n$ increases the power in both cases. However, as known from (4), the power under the normal design prior is bounded by $1-\Phi\{(\theta_{0}+\mu-2\mu_{d})/(2\tau_{d})\}=77.3\%$ while the power under the point design prior can be increased up to $100\%$. Finally, the curves cross at 50%, the intuition being that with the normal design prior there is more uncertainty, and hence the power is always closer to $50\%$.
@@
see Appendix B for details. Unlike the power function based on the Bayes factor with point analysis prior (3), the power function based on the Bayes factor with normal analysis prior (5) can be increased arbitrarily close to one by increasing the sample size $n$ (see Appendix C). That is, we have that

### v2

@@ Abstract
Determining an appropriate sample size is a critical element of study design, and the method used to determine it should be consistent with the planned analysis. When the planned analysis involves Bayes factor hypothesis testing, the sample size is usually desired to ensure a sufficiently high probability of obtaining a Bayes factor indicating compelling evidence for a hypothesis, given that the hypothesis is true. In practice, Bayes factor sample size determination is typically performed using computationally intensive Monte Carlo simulation. Here, we summarize alternative approaches that enable sample size determination without simulation. We show how, under approximate normality assumptions, sample sizes can be determined numerically, and provide the R package bfpwr for this purpose. Additionally, we identify conditions under which sample sizes can even be determined in closed-form, resulting in novel, easy-to-use formulas that also help foster intuition, enable asymptotic analysis, and can also be used for hybrid Bayesian/likelihoodist design. Furthermore, we show how power and sample size can be computed without simulation for more complex analysis priors, such as Jeffreys-Zellner-Siow priors or non-local normal moment priors. Case studies from medicine and psychology illustrate how researchers can use our methods to design informative yet cost-efficient studies.
Keywords: Bayesian hypothesis testing, design prior, evidence, likelihood ratio, study design
@@ 1 Introduction
A key aspect of study design is determining an appropriate sample size. Choosing a sample size that is too small may lead to inconclusive study results, while choosing a sample size that is too large may be unethical (e.g., for animal studies) or waste samples which could be of better use in other studies. Whether or not a certain sample size can ensure sufficiently conclusive results depends on the planned analysis. Therefore, the sample size calculation should be aligned with the goals of the analysis (Anderson and Kelley, , 2024), or in other words: ‘As ye shall analyse is as ye shall design(Julious, , 2023, p. 179).
@@
where $z_{q}$ is the $q\times 100\%$ quantile of the standard normal distribution, $\alpha$ is the level of the test, $1-\beta$ is the desired power, $\mu$ is the assumed mean difference, $\theta_{0}$ is the mean difference under the null hypothesis, and $\sigma^{2}$ is the variance of one observation (Matthews, , 2006, p. 34). There exist various refinements of (1), such as, adaptations to unequal randomization, special study designs (e.g., cross-over studies), or other data types (e.g., binary data), see, for example, Kieser, (2020) or Julious, (2023). For many analysis methods, however, no closed-form formula exist and iterative or simulation methods have to be used. Nevertheless, while typically only being an approximation, the formula (1) enables ‘quick-and-dirty’ calculations that are often accurate enough for practical purposes. It also helps fostering intuition and is therefore useful, for example, in teaching of statistics. Finally, the formula is helpful for theoretical derivations such as ‘how much do we need to increase the sample size if we reduce $\alpha=0.05$ down to $\alpha=0.005$?’ (Benjamin et al., , 2017).
An alternative to frequentist hypothesis testing is Bayesian hypothesis testing. There are different flavors of Bayesian hypothesis testing, one of the most popular being the approach centered around the Bayes factor, which is the data-based updating factor of the prior to posterior odds of two competing hypotheses. Bayes factor approaches were pioneered by Jeffreys, (1939) and are now in use in various scientific domains such as medicine (Goodman, , 1999), psychology (Morey et al., , 2016; Heck et al., , 2023), or physics (Trotta, , 2008). Bayes factor tests are conceptually different from frequentist tests in several ways. For example, they can quantify evidence in favor of a null hypothesis or they can incorporate external information via a prior distribution. For an overview of Bayes factors see e.g., Kass and Raftery, (1995); Held and Ott, (2018).
@@
The goal of this paper is therefore to investigate whether, under approximate normality assumptions similar to those underlying the formula (1), a sample size formula can be derived for a planned Bayes factor analysis. As we will show, the answer is affirmative under certain assumptions about the analysis prior (the prior distribution for the parameter used in the analysis) and the design prior (the prior distribution for the parameter used in the design). A distinction must be made between point priors and normal priors, both in the design and in the analysis. Point analysis priors lead to Bayes factors reducing to likelihood ratios, the analysis thereby corresponding with a frequentist ‘likelihoodist’ analysis (Edwards, , 1971; Royall, , 1997; Blume, , 2002; Strug, , 2018; Cahusac, , 2020), while point design priors lead to ‘conditional power’, which corresponds to traditional frequentist power. In contrast, normal analysis and design priors can account for parameter uncertainty, producing Bayes factors that differ from likelihood ratios and ‘predictive power’ that differs from frequentist power (see e.g., O'Hagan et al., , 2005; Micheloud and Held, , 2022; Grieve, , 2022, for conditional/predictive power related to posterior tail probability analyses).
Based on the point/normal prior distinction, we find that closed-form sample sizes are available for Bayes factors with point analysis priors (i.e., likelihood ratios) along with point or normal design priors, and for Bayes factors with local normal analysis and design priors (i.e., normal priors centered on the null value). Table 1 provides an overview of our results. In addition to our novel formulas, we summarize sample size determination for Bayes factors with normal priors based on numerical root-finding (which has been done before, e.g., in Weiss, , 1997; Pawel et al., , 2023), and show how it can be extended to more advanced prior distributions, such as normal moment priors and Jeffreys-Zellner-Siow priors. While root-finding approaches also do not produce closed-form sample sizes, computations are deterministic and usually faster than with simulation approaches. To facilitate reuse of our results, all methods are made available through our R package bfpwr.
@@
This paper is organized as follows: We begin by defining the type of Bayes factor underlying our sample size calculations (Section 2), followed by deriving its distribution when new data are generated under various design priors (Section 3). Combining these results, we derive several formulas for the sample size under different constellations of design and analysis priors (Section 4). Examples from medicine and psychology then illustrate how our formulas can be used by researchers in practice (Section 5). Section 6 illustrates possible extensions of our framework to other popular types of analysis priors, such as informed $t$ priors (Gronau et al., , 2020) and non-local normal moment priors (Johnson and Rossell, , 2010). The paper ends with a closing discussion of our results and final remarks on limitations and extensions (Section 7). Our R package bfpwr that implements the developed methods is illustrated in Appendix A.
2 Bayes factor analysis
@@
To derive a sample size formula, we must first clarify how the future data will be analyzed. Denote by $\hat{\theta}$ the estimate of an unknown parameter $\theta$ that will result from the statistical analysis of a future data set with effective sample size $n$. Suppose that the estimate’s standard error is of the form $\sigma_{\scriptscriptstyle\hat{\theta}}/\sqrt{n}$ where $\sigma_{\scriptscriptstyle\hat{\theta}}^{2}$ is the variance of one effective observation. For example, if the data are normally distributed with known variance $\sigma^{2}$ and the parameter of interest is their mean $\theta$ estimated with the sample mean $\hat{\theta}$ (the ‘one-sample’ case), we have that $\mathrm{Var}(\hat{\theta})=\sigma^{2}/n$ so the unit variance is $\sigma_{\scriptscriptstyle\hat{\theta}}^{2}=\sigma^{2}$ and the effective sample size $n$ is the number of observations. If there were another group of $n$ normally distributed observations with a potentially different mean but the same known variance $\sigma^{2}$, and we were interested in the mean difference $\theta$ estimated with the empirical mean difference $\hat{\theta}$ (the ‘two-sample’ case), we would have that $\mathrm{Var}(\hat{\theta})=(2\sigma^{2})/n$ so the unit variance would be $\sigma_{\scriptscriptstyle\hat{\theta}}^{2}=2\sigma^{2}$ and the effective sample size $n$ would be the number of observations per group. For many commonly used estimators, it is reasonable to assume that the estimate (but not necessarily the underlying data) is approximately normally distributed around $\theta$ with variance equal to the squared standard error, i.e., $\hat{\theta}\mid\theta\sim\mathrm{N}(\theta,\sigma^{2}_{\scriptscriptstyle\hat{\theta}}/n)$. Table 2 shows common types of parameter estimates and the resulting interpretation of the effective sample size $n$ and the unit variance $\sigma^{2}_{\scriptscriptstyle\hat{\theta}}$. It is important to note that these are only approximations and they may be inadequate in certain situations, such as small sample sizes or when the assumed variance $\sigma^{2}$ is strongly misspecified (Spiegelhalter et al., , 2004). The frequentist sample size formula (1) can also be cast in this framework; it assumes continuous outcome data and a mean difference parameter (second row in Table 2). As such, the formula could be generalized to other settings by replacing $2\sigma^{2}$ in the numerator with other unit variances from Table 2, which would in turn change the interpretation of $n$ and $\hat{\theta}$.
@@
Table 2: Different types of parameter estimates $\hat{\theta}$ with approximate variance $\mathrm{Var}(\hat{\theta})=\sigma^{2}_{\scriptscriptstyle\hat{\theta}}/n$ and corresponding interpretation of sample size $n$ and unit variance $\sigma^{2}_{\scriptscriptstyle\hat{\theta}}$ (adapted from Chapter 2.4 in Spiegelhalter et al., , 2004 and Chapter 1 in Grieve, , 2022). The variance of one continuous outcome observation is denoted by $\sigma^{2}$ and assumed to be known. Parameter estimates based on two groups assume an equal number of observations per group.
@@
Binary Arcsine square root difference Sample size per group 1/2
@@
A Bayes factor less than one ($\text{BF}_{01}<1$) indicates evidence for the alternative hypothesis $H_{1}$, while a Bayes factor greater than one ($\text{BF}_{01}>1$) indicates evidence for the null hypothesis $H_{0}$. The larger the deviation of the Bayes factor from one, the stronger the evidence. Conventional thresholds for substantial and strong evidence for the alternative (null) hypothesis are $1/3$ and $1/10$ ($3$ and $10$), respectively. A Bayes factor 1/$31/3<\text{BF}_{01}<$ 3 is typically interpreted as absence of evidence for either hypothesis, calling for more data to discern the more appropriate hypothesis (Jeffreys, , 1939; Held and Ott, , 2018).
The Bayes factor (2) is implemented in our package in the function bf01. It has already appeared in the literature in one form or another (e.g., in Weiss, , 1997; De Santis, , 2004; Spiegelhalter et al., , 2004; Dienes, , 2014; Bartoš and Wagenmakers, , 2023), with perhaps the first proposal of a Bayes factor based on an approximately normally distributed parameter estimate and its standard error dating back to Jeffreys, (1936), see also Wagenmakers, (2022) for some historical notes on Jeffreys’ approach. The Bayes factor (2) may be thought of as a ‘Bayesian $z$-test’, that is, a test of a normal mean based on an asymptotically normal statistic assuming that the variance of the statistic is known. Of course, the latter assumption is not true in most applications, but it makes the test widely applicable and is, for practical purposes, often close enough to Bayes factors based on the exact distribution of the data, which may or may not be available. Finally, the Bayes factor (2) can also be used with parameter estimates where a standard error is available but not of the form $\sigma_{\scriptscriptstyle\hat{\theta}}/\sqrt{n}$, e.g., a parameter estimate from a generalized linear model where the estimate is adjusted for covariates and the standard error is obtained numerically. In this case, $\sigma_{\scriptscriptstyle\hat{\theta}}/\sqrt{n}$ in (2) can be replaced by the observed standard error. However, as we will show now, assuming such a particular dependence on the sample size $n$ allows us to perform closed-form power and sample size calculations under certain additional assumptions.
@@ 3 Distribution and power function of the Bayes factor
Suppose now that we are interested in finding compelling evidence – either in favor of the alternative $H_{1}$ over the null hypothesis $H_{0}$ with a Bayes factor (2) smaller than some threshold $k<1$ (e.g., $k=1/3$ or $k=1/10$) or in favor of $H_{0}$ over $H_{1}$ with a Bayes factor greater than $k>1$ (e.g., $k=3$ or $k=10$). To determine a sample size that ensures compelling evidence with a desired probability we need to know the distribution of the Bayes factor (2) for a given sample size.
Assume a so-called ‘design prior’ for the parameter $\theta$ that is used in the design of the study (O'Hagan and Stevens, , 2001; O'Hagan et al., , 2005). This prior should represent the state of knowledge and uncertainty about $\theta$ at the design stage and does not necessarily have to correspond to the ‘analysis prior’ $\theta\mid H_{1}\sim\mathrm{N}(\mu,\tau^{2})$ used for the Bayes factor (2). In fact, the analysis prior is often set to a certain ‘default’ or ‘objective’ prior that is conventionally used in the field. Here, we will focus on normal design priors $\theta\sim\mathrm{N}(\mu_{d},\tau^{2}_{d})$, as they are flexible enough to specify varying degrees of uncertainty about the parameter, and at the same time mathematically convenient for obtaining closed-form solutions for power and, in some cases, sample size. Point priors, and as such classical sample size determination with an ‘assumed parameter’, then represent a special case of the normal prior where the standard deviation becomes infinitesimally small ($\tau_{d}=0$). In addition, specifying a point design prior at the null value ($\mu_{d}=\theta_{0}$ and $\tau_{d}=0$) allows us to compute the probability of finding compelling evidence for the true null hypothesis $H_{0}$, as well as the probability of finding misleading evidence for the alternative when the null hypothesis is true (the ‘type I error rate’). Presenting the latter probability can be useful for demonstrating that the chosen design is appropriately ‘calibrated’ (Dawid, , 1982; Rubin, , 1984; Little, , 2006; Grieve, , 2016).
@@
$Z=\frac{1}{\sqrt{\tau^{2}_{d}+\sigma^{2}_{\scriptscriptstyle\hat{\theta}}/n}}\left\{\frac{\sigma^{2}_{\scriptscriptstyle\hat{\theta}}\log k}{n(\theta_{0}-\mu)}+\frac{\theta_{0}+\mu}{2}-\mu_{d}\right\},$ (4)
see Appendix B for details. We may also want to compute the probability that the Bayes factor is greater than $k$ (e.g., to determine the probability of compelling evidence for a true null hypothesis), which is simply one minus the probability (3).
In standard frequentist sample size determination, the power can typically be increased arbitrarily close to one by increasing the sample size. However, with the Bayes factor based on a point analysis prior, depending on the assumed design prior, one may not be able to approach a power of one with the power function (3) by increasing the sample size $n$. That is, the limiting power value is given by
@@
with $Z_{\mathrm{lim}}=(\theta_{0}+\mu-2\mu_{d})/(2\tau_{d})$, the limit of (4) for increasing $n$. When the design prior is also a point prior ($\tau_{d}=0$), $Z_{\mathrm{lim}}$ diverges and the limiting power (5) approaches one or zero, depending on whether the location of the design prior $\mu_{d}$ is closer to the alternative $\mu$ or to the null $\theta_{0}$. In case it is just in between the two ($\mu_{d}=(\theta_{0}+\mu)/2$), the limiting power approaches a half. On the other hand, for a normal design prior ($\tau_{d}>0$), the limiting power is bounded by a value in between (and not including) zero and one given by (5). The intuition behind these results is that for a normal design prior, there is always parameter uncertainty, even if the sample size becomes arbitrarily large, while for point design priors, the parameter uncertainty can be arbitrarily reduced by increasing the sample size. This parallels similar results on bounds for hybrid Bayesian/frequentist power (Spiegelhalter et al., , 2004; Micheloud and Held, , 2022; Grieve, , 2022).
@@
The left plot in Figure 1 shows examples of two power functions related to the Bayes factor for a standardized mean difference parameter with point analysis prior. The power is computed under either a point (red curve) or a normal design prior (blue curve). We see that increasing the sample size $n$ increases the power in both cases. However, as known from (5), the power under the normal design prior is bounded by $1-\Phi(Z_{\mathrm{lim}})=77.3\%$ while the power under the point design prior can be increased up to $100\%$. Finally, the curves cross at 50%, the intuition being that with the normal design prior there is more uncertainty, and hence the power is always closer to $50\%$.
@@
see Appendix B for details. Again, to compute the probability of a Bayes factor in favor of the null hypothesis, we have to take one minus the probability (6) along with a level $k>1$.
Unlike the power function based on the Bayes factor with point analysis prior (3), the power function based on the Bayes factor with normal analysis prior (6) can be increased arbitrarily close to one by increasing the sample size $n$ (see Appendix C). That is, we have